# Calibration Attention: Instance-wise Temperature Scaling for Vision Transformers

### v1

@@ 3.2 Main calibration results (ECE)
3.3 Fine-grained reliability: AdaECE, smECE and Classwise-ECE
The classical Expected Calibration Error (ECE) aggregates reliability with a fixed histogram; empty or highly-skewed bins may therefore hide serious local mis-calibration. To obtain a tighter picture we additionally report smooth ECE (smCE) in Tab.4, Adaptive ECE (AdaECE) in Tab. 5 and Classwise-ECE in Tab. 6. Smooth ECE. On CIFAR-100, CalAttn reduces smCE from 2.17 % to 1.45 % on ViT224 (a $33$ % drop) and from 1.75 % to 1.17 % on DeiTSmall ($32$ %). The best overall score is obtained by SwinSmall+CalAttn with 1.64 %. Adaptive ECE. On CIFAR-10, ViT224+CalAttn reaches 1.10 %, halving the tuned baseline (49 % reduction). Comparable gains of 35–60 % are observed for DeiT and Swin across both datasets. Classwise-ECE. CalAttn converges to 0.26–0.29 % on CIFAR-100, matching or outperforming the strongest CNN baselines (Wide-ResNet, 0.20 % with specialised losses) while adding less than 0.1 % additional parameters (Table 7).
3.4 Reliability diagrams and Robustness on OoD Data Shift
Visual evidence of calibration gains.
Figure 3 and Figure 4 complement the quantitative Tables 34. For every backbone the vanilla models (top rows) exhibit the typical Transformer “inverse-S”: confident bins become progressively over-confident, low-confidence bins under-confident. Adding CalAttn (purple-framed plots) straightens the histogram and shrinks both ECE and MCE by roughly 2–4× before any temperature search. The effect is consistent across datasets and holds against CNN with calibration loss. Robustness on OoD data shift is provided in Appendix I.
Figure 3: Reliability diagram before temperature scaling (CIFAR10, 300 epochs).
Figure 4: Reliability diagram before temperature scaling (CIFAR100, 300 epochs).
3.5 Ablation on CalAttn Input Features and Head Type
We investigate how the choice of feature vector fed to CalAttn influences calibration, keeping all other settings (e.g., MLP depth, learning rate, loss) fixed. CLS (original). CalAttn receives the final [CLS] token ($B\times D$). Patch-mean. The [CLS] token is replaced with the mean of all patch tokens, providing a global spatial summary. This simple change improves CIFAR-100 ECE from $6.87\%$ to $4.38\%$ on DeiT-S. Concat. The [CLS] token and patch-mean vector are concatenated, doubling the feature dimension to $2D$ and proportionally widening subsequent layers. Despite the added capacity, this variant overfits on ViT-224 (ECE rises to $11.56\%$) and is retained only as a diagnostic baseline.
Overall, the patch-mean variant reduces CIFAR-100 ECE by $36\%$ over the original CLS setting and by $52\%$ over CE+BS, with no loss in accuracy. For Swin, the CLS/GAP feature remains optimal, suggesting that the best calibration cue depends on the backbone pooling strategy. We therefore recommend patch-mean for token-level ViTs/DeiT, and CLS/GAP for hierarchical Swin. Full numerical results are provided in Appendix J. We also performed ablation on different head type: by scalar and Dirichlet $\alpha$-Head in Table 13.
3.6 Results on ImageNet-1K
We further evaluate CalAttn on ImageNet-1K, fine-tuning for 350 epochs, and compare against recent calibration baselines Luo et al. [2025]. Using Swin-S as the backbone, CalAttn reduces ECE from $4.95\%$ (CE+BS) to $1.25\%$—a 75% relative drop—and lowers high-confidence error (AECE) by $20\%$, while maintaining top-1 accuracy within $0.1$ pp of the strongest baseline. Full results are provided in Appendix J.1. We also compare CalAttn against recent SATS Joy et al. [2023] under the standard post-hoc temperature scaling protocol in Appendix J.2.
4 Discussion
4.1 Interpreting the gains of CalAttn
From hidden “thermometer” to usable temperature. Tables 36 show that CalAttn cuts every calibration metric by $40\!-\!55\%$ before any post-hoc search. The reliability diagrams in Fig. 3 and Fig. 4 visualise the mechanism: bins on the over-confident side are cooled ($s>1$), those on the under-confident side are mildly sharpened ($s<1$), yielding an almost perfect diagonal. Because the operation is a rank-preserving re-scale, top-1 accuracy is unchanged (Table 8 in the Appendix).
Heteroscedastic view. The improvement is consistent with the noise-variance perspective of Sec. 2.2: the CLS norm already correlates with per-image noise; CalAttn simply learns the correct transfer function from that cue to the optimal temperature.
4.2 Transformer vs. CNN calibration
After optimal scalar $T^{\star}$ the supposed “calibration premium” of Transformers largely disappears: on CIFAR-100, tuned ViT-224 sits at $3.3\%$ ECE, on par with a post-hoc–scaled Wide-ResNet-26-10 ($2.8\%$). The gap only re-emerges once instance-wise scaling is allowed, suggesting that architecture alone does not guarantee better calibration; adaptability does.
4.3 Temporal behaviour of CalAttn
Figure 5: Dynamics of the temperature head on ViT-224 (CIFAR-10). Top: mean scale $s(z)$ (red, right axis) and coefficient of variation $\mathrm{CV}_{s}$ (blue, left axis). Grey dashed line marks the best global temperature $T^{\star}$. Shaded bands denote cooling ($s>1$) and sharpening ($s<1$). Bottom: kernel-density estimates of $s(x)$ and $\lVert z_{\mathrm{CLS}}\rVert_{2}$ at three epochs; legends report the mean $\mu$ and CV.
Figure 5 traces the per-sample scale learned by CalAttn over 350 epochs (cosine schedule, LR $\times 0.1$ at epoch 150). Three distinct phases emerge:
Phase I – global alignment (epochs 0–30). The mean scale collapses from 40 to 4 while the coefficient of variation (CV) plummets from $0.35$ to $0.05$. The head therefore behaves like classical temperature scaling, quickly discovering a value that closely matches the post-hoc optimum $T^{\star}$.
Phase II – selective sharpening (epochs 30–150). After a stable plateau the LR drop triggers a second descent of the mean: $s(z)$ crosses the neutral line s=1s\!=\!1 and enters the sharpening regime (green band). The CV bottoms out ($\approx\!0.02$), indicating that most images still share a similar scale—CalAttn has not yet begun to exploit heteroscedasticity.
Phase III – heteroscedastic adaptation (epochs 150–350). Once the backbone has largely converged, the CV rises steadily to $\approx\!0.18$: CalAttn now assigns distinct temperatures, cooling over-confident instances (s>1s\!>\!1) and sharpening under-confident ones (s<1s\!<\!1). Fig.5 confirm that the scale distribution widens while the CLS-norm distribution remains nearly stationary; the head therefore extracts additional calibration gain without perturbing the representation.
Take-away. CalAttn first replicates the best global $T^{\star}$, then learns instance-wise deviations that track sample difficulty. The late-epoch rise of $\operatorname{CV}_{s}$ is thus desirable: it quantifies the head’s heteroscedastic expressiveness and correlates with the extra ECE reduction observed after epoch 150 (cf. Table 3). Crucially, the mean settling below one shows that its logits are not merely over-confident; some classes benefit from sharpening, that global temperature scaling cannot capture.
5 Related Work
Vision Transformers and their descendants. The seminal Vision Transformer (ViT) showed that a pure Transformer encoder—long dominant in NLP—can rival strong CNN baselines when trained on sufficiently large image corpora [Dosovitskiy et al., 2021]. Its heavy data requirements were soon alleviated by stronger regularisation and knowledge–distillation; DeiT reaches ResNet-50 accuracy on ImageNet using standard hardware and no extra data [Touvron et al., 2021]. Parallel lines of work addressed the quadratic cost of full self-attention. The Swin Transformer introduces shifted window attention and a hierarchical pyramid, becoming the de-facto Transformer counterpart of FPNs for dense prediction [Liu et al., 2021]. Further refinements explore tokenisation [Yuan et al., 2021], lightweight mixing [Chen et al., 2021], and positional priors [Dai et al., 2021]. Across this rapidly expanding family, however, the inference pipeline still terminates in a fixed soft-max layer; trustworthy confidence therefore hinges on post-hoc calibration.
Calibration for transformer classifiers. Classical calibration traces back to Platt scaling [Platt et al., 1999]; its modern successor, global temperature scaling (TS), remains the community’s work-horse [Guo et al., 2017]. Early studies focussed on CNNs, but the question of Transformer calibration has gained momentum only recently. Minderer et al. [2021] report that ViTs are often better calibrated than CNNs, yet still benefit from a tuned $T^{\star}$ and reverses under distribution shift. Several losses have been proposed to side–step the global-temperature bottleneck. Adaptive focal variants increase the weight of poorly calibrated samples [Mukhoti et al., 2020a]; label smoothing can add a mild regularising signal [Müller et al., 2019]. Dual Focal Loss (DFL) enlarges the gap between the top two logits to reduce over-confidence [Tao et al., 2023]; Focal Calibration Loss (FCL) blends focal loss with a differentiable ECE surrogate [Liang et al., 2024b]. Although effective, these objectives rely on fixed hyper-parameters and still apply the same scale to every image. Ensemble methods are also raised to improve a model’s calibration and robustness [Gal and Ghahramani, 2016, Lakshminarayanan et al., 2017, Ovadia et al., 2019, Wen et al., 2020]. CTKD [Li et al., 2023] and MKD [Liu et al., 2022a] schedule a dynamic temperature during teacher-student distillation to balance hard and soft targets. CSM [Luo et al., 2025] is a diffusion-based re-annotation strategy for mixed-label samples to calibrate.
Gap addressed by Calibration Attention. Existing methods either (i) apply a single global temperature after training [Guo et al., 2017, Kumar et al., 2018, Mukhoti et al., 2020a, Joy et al., 2023] or (ii) prescribe static loss hyper-parameters [Tao et al., 2023, Liang et al., 2024b]. Neither class explicitly models the heteroscedastic uncertainty present across samples. Our CalAttn head learns an instance-wise temperature from the hidden CLS representation, integrates seamlessly with any ViT-style backbone, and is trained jointly with the task loss—bridging the remaining gap between predictive accuracy and reliable uncertainty. CalAttn is designed for vision transformers (ViT, DeiT, Swin) and relies on the presence of a global token (e.g.[CLS]) or equivalent GAP feature.
6 Conclusion
Post-hoc temperature–scaling remains the dominant fix for neural-network mis-calibration, yet its single global parameter cannot capture the wide variability in sample difficulty. CalAttn is designed for vision transformers (ViT, DeiT, Swin) and relies on the presence of a global token (e.g. [CLS]) or equivalent GAP feature. We have not yet explored extensions to CNNs, hybrid CNN-ViT models, or non-vision modalities. Adapting CalAttn to architectures without an explicit global token—by, for example, learning spatial pooling schemes—or to text and multimodal transformers is promising future work.
Appendix A Notation table for the Calibration-Attention
All notations in our work define in Table 3.
Table 3: Notation and abbreviations used throughout the paper. Shapes refer to vision experiments.
Symbol / op. Meaning Shape / type
$B$ Mini–batch size
$C$ Number of classes
NN Number of image patches
$L$ Transformer encoder depth
$d$ Token / CLS embedding dimension e.g. 384, 768
$H,W$ Input image height / width e.g. 32, 224
$\mathbf{x}$ Input image tensor $(B,3,H,W)$
patch_embed($\cdot$) Linear projection of image patches
$\text{CLS}]$ Learned classification token $(1,d)$
$\mathbf{z}_{\text{CLS}}\;(=\mathbf{z})$ Final CLS embedding (after encoder & LN) $(B,d)$
$W_{\text{cls}},\,b_{\text{cls}}$ Weights / bias of classifier head $d{\times}C,\;C$
@@ Appendix A Notation table for the Calibration-Attention
$h$ Hidden width of Calibration MLP default 128
$\mathbf{h}$ Hidden activation $\mathrm{GELU}(\mathbf{z}W_{c}^{(1)})$ $(B,h)$
Calibration Head Two-layer MLP: $\mathbf{z}\!\xrightarrow{\text{FC}(d\!\to\!h)}\!\xrightarrow{\text{GELU}}\!\xrightarrow{\text{FC}(h\!\to\!1)}\!\xrightarrow{\text{Softplus}}\!s(\mathbf{z})$
$s(\mathbf{z})$ Per-sample temperature $(B,1)$
$\varepsilon$ Numerical safety term ($10^{-6}$) scalar
$\tilde{\boldsymbol{\ell}}$ Scaled logits $\boldsymbol{\ell}/(s(\mathbf{z})+\varepsilon)$ $(B,C)$
@@
$\mathbf{e}_{y}$ One-hot ground-truth label $(B,C)$
$\mathcal{L}_{\mathrm{CE}}$ Cross-entropy loss scalar
$\mathcal{L}_{\mathrm{Brier}}$ Brier penalty $\|\hat{\mathbf{y}}-\mathbf{e}_{y}\|_{2}^{2}$ scalar
$\lambda$ Weight on Brier term (default 0.1) scalar
ECE, AdaECE, smCE, Classwise-ECE Calibration metrics
Operators / special symbols
$\boxplus$ Element-wise addition (broadcast) with positional encoding binary op.
$\|\cdot\|_{2}$ Euclidean (L2) norm scalar
$\operatorname{LN}(\cdot)$ Layer Normalisation function
$\sigma_{i}$ $i$-th component of the soft-max scalar prob.
$\operatorname{GELU}$, $\operatorname{Softplus}$ Non-linearities
Appendix B Calibration Metrics
A classifier is perfectly calibrated when the predictive confidence matches the empirical accuracy, i.e. $\Pr\!\bigl{(}y=\hat{y}\,\big{|}\,\hat{p}=p\bigr{)}=p$ for every $p\in[0,1]$. All metrics below measure the deviation from this ideal diagonal.
Histogram-based metrics.
Let $\hat{p}_{i}=\max_{c}\hat{y}_{ic}$ be the top-class confidence for sample $i$ and $\mathbbm{1}(\cdot)$ the indicator function. Partition the unit interval into $M$ bins and define
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\!\sum_{i\in B_{m}}\!\mathbbm{1}(\hat{y}_{i}=y_{i}),\qquad\text{conf}(B_{m})=\frac{1}{|B_{m}|}\!\sum_{i\in B_{m}}\!\hat{p}_{i}.$ (7)
With $N=\sum_{m}|B_{m}|$ the classic metrics become
$\operatorname{ECE}$ $\,=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\bigl{|}\text{acc}(B_{m})-\text{conf}(B_{m})\bigr{|},$ (8)
$\operatorname{MCE}$ $\,=\max_{m}\bigl{|}\text{acc}(B_{m})-\text{conf}(B_{m})\bigr{|},$ (9)
$\operatorname{AdaECE}$ $\,=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\bigl{|}\text{acc}(B_{m})-\text{conf}(B_{m})\bigr{|}\quad\text{(adaptive binning)},$ (10)
$\operatorname{ClassECE}$ $\,=\frac{1}{K}\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{|B_{km}|}{N_{k}}\bigl{|}\text{acc}(B_{km})-\text{conf}(B_{km})\bigr{|}.$ (11)
Smooth ECE (smECE).
Hard bin edges can mask fine-grained mis-calibration. Błasiok and Nakkiran [2023] replace binning with kernel smoothing. Using a Gaussian kernel $K_{h}(t)=\exp\!\bigl{(}-t^{2}/2h^{2}\bigr{)}$ and bandwidth $h$:
$\widehat{\text{acc}}(\hat{p}_{i})=\frac{\sum_{j}K_{h}(\hat{p}_{i}-\hat{p}_{j})\,\mathbbm{1}(\hat{y}_{j}=y_{j})}{\sum_{j}K_{h}(\hat{p}_{i}-\hat{p}_{j})},\qquad\operatorname{smECE}=\frac{1}{N}\sum_{i=1}^{N}\bigl{|}\widehat{\text{acc}}(\hat{p}_{i})-\hat{p}_{i}\bigr{|}.$ (12)
The kernel view removes bin-edge artefacts and yields a differentiable objective, which underpins recent calibration losses such as FCLLiang et al. [2024b] and the proposed CalAttn head. [^4] We follow Błasiok and Nakkiran [2023] and fix $h=0.05$ for all experiments.
Interpretation in this work.
ECE, MCE, AdaECE and Classwise-ECE are intuitive and readily visualised via reliability diagrams, while smECE offers a high-resolution, bin-free perspective. We report all five scores (Sec. 3) and show that CalAttn lowers each of them— including the stringent smECE—demonstrating genuine, sample-wise calibration improvements beyond coarse histogram effects.
Appendix C Architectural Differences
We show differences between the vanilla Vision Transformer and our proposed Calibration-Attention variant in Table 4.
Table 4: Architectural differences between the vanilla Vision Transformer and our proposed Calibration-Attention variant.
Component Baseline ViT ViT+Calibration Attention
Extra module Calibration Head — a two-layer MLP attached to the CLS embedding
Output of the module Strictly positive scale $s(\mathbf{z})$ (initially $\approx 1.0$)
@@ Appendix C Architectural Differences
Loss function Cross-entropy only Cross-entropy +\;+\; small Brier term (differentiable calibration surrogate)
Post-hoc temperature search Required Not needed — network learns $s(\mathbf{z})$ end-to-end
Extra parameters 0 $0.1\%$ of total (two small fully-connected layers)
Appendix D Vision Transformer + Calibration Attention: Forward Steps
1. Patch embedding.
$\mathbf{X}_{\text{patch}}=\text{reshape}\!\bigl{(}\!\text{Conv}_{3\!\to\!D}^{P,P}(\mathbf{x}),\,N,\,D\bigr{)}\quad\in\mathbb{R}^{N\times D}$
2. Class token and positions.
$\mathbf{z}^{(0)}=\bigl{[}\;\mathbf{z}_{\text{cls}}\;\;\mathbf{X}_{\text{patch}}\bigr{]}\boxplus\mathbf{E}_{\text{pos}}\;\in\;\mathbb{R}^{(N+1)\times D}$
3. Encoder blocks  (for $\ell=1..L$).
Each block contains multi-head self-attention (MSA) followed by a feed-forward network (MLP) with residual connections and LayerNorm:
$\mathbf{u}^{(\ell)}$ $=\text{LN}\!\bigl{(}\mathbf{z}^{(\ell-1)}\bigr{)},$
$\mathbf{a}^{(\ell)}$ $=\text{MSA}\!\bigl{(}\mathbf{u}^{(\ell)}\bigr{)}=\bigoplus_{k=1}^{h}\underbrace{\text{softmax}\!\Bigl{(}\tfrac{\mathbf{Q}_{k}\mathbf{K}_{k}^{\top}}{\sqrt{D/h}}\Bigr{)}\mathbf{V}_{k}}_{\text{head }k}\mathbf{W}_{o},$
$\mathbf{z}^{(\ell)}$ $=\mathbf{z}^{(\ell-1)}+\mathbf{a}^{(\ell)},$
$\mathbf{v}^{(\ell)}$ $=\text{LN}\!\bigl{(}\mathbf{z}^{(\ell)}\bigr{)},$
$\mathbf{m}^{(\ell)}$ $=\text{MLP}\!\bigl{(}\mathbf{v}^{(\ell)}\bigr{)}=\sigma\!\bigl{(}\mathbf{v}^{(\ell)}\mathbf{W}_{1}\bigr{)}\mathbf{W}_{2},$
$\mathbf{z}^{(\ell)}$ $=\mathbf{z}^{(\ell)}+\mathbf{m}^{(\ell)}.$
4. Calibration head (instance–wise temperature).
$\mathbf{h}$ $=\text{GELU}\!\bigl{(}\mathbf{z}^{(L)}_{0}\mathbf{W}_{c}^{(1)}\bigr{)}\quad\in\mathbb{R}^{1\times H},$
$s$ $=\text{softplus}\!\bigl{(}\mathbf{h}\mathbf{W}_{c}^{(2)}+b_{c}\bigr{)}+\varepsilon\quad\in\mathbb{R}_{>0},$
$\boldsymbol{\ell}$ $=\frac{\mathbf{z}^{(L)}_{0}\mathbf{W}_{\text{cls}}}{s}\quad\in\mathbb{R}^{C},$
$\hat{\mathbf{y}}$ $=\text{softmax}\!\bigl{(}\boldsymbol{\ell}\bigr{)}.$
5. Loss.
$\mathcal{L}=-\log\hat{y}_{y}\;+\;\lambda\,\|\hat{\mathbf{y}}-\mathbf{e}_{y}\|_{2}^{2}\quad(\lambda=0.1).$
Appendix E Proof that CalAttn converges to the optimal temperature
[…]
$\mathcal{L}=-\log\hat{y}_{y}+\lambda\!\left\|\hat{\mathbf{y}}-\mathbf{e}_{y}\right\|_{2}^{2}\;.$
The derivative w.r.t. $s$ is [^5] See Guo et al. [2017] for the CE part; the Brier term is similar.
$\frac{\partial\mathcal{L}}{\partial s}=\bigl{(}\hat{y}_{y}-1\bigr{)}\frac{\ell_{y}-\sum_{j}\hat{y}_{j}\ell_{j}}{s}+2\lambda\sum_{c}\bigl{(}\hat{y}_{c}-\mathbf{e}_{yc}\bigr{)}\hat{y}_{c}(1-\hat{y}_{c})\frac{\ell_{c}-\sum_{j}\hat{y}_{j}\ell_{j}}{s}.$
Setting $\partial\mathcal{L}/\partial s=0$ and rearranging yields
$s^{\star}=\frac{\ell_{y}-\sum_{j}\hat{y}_{j}\ell_{j}}{\ell_{y}^{\mathrm{opt}}-\sum_{j}\hat{y}_{j}^{\mathrm{opt}}\ell_{j}},$
where “opt” denotes the soft-max that exactly matches empirical accuracy, $\hat{y}_{y}^{\mathrm{opt}}=\mathbf{1}(\hat{y}=y)$. Hence $s^{\star}$ equals the temperature that aligns confidence with correctness for that sample—i.e. the per-instance optimum of temperature scaling.
Appendix F AdaECE and Classwise-ECE
We extend the metrics to AdaECE and Classwise-ECE in Tables 5 and 6.
Table 5: $\downarrow$ AdaECE (%) for methods both before and after applying temperature scaling.
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
@@ Appendix F AdaECE and Classwise-ECE
Table 6: $\downarrow$ Classwise-ECE (%) for methods both before and after applying temperature scaling.
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T

### v2

@@ Appendix I Sensitivity to the CE+Brier Trade-off $\lambda$ (CIFAR-100, DeiT-S)
$\lambda$ ECE $\downarrow$ AdaECE $\downarrow$ ClassECE $\downarrow$ smECE $\downarrow$
0.1 2.39 2.44 0.26 2.44
0.2 2.48 2.38 0.27 2.47
0.3 2.50 2.39 0.27 2.48
0.4 2.48 2.38 0.28 2.47
0.5–1.0 2.48 2.38 0.27 2.47
@@
$\lambda$ ECE $\downarrow$ AdaECE $\downarrow$ ClassECE $\downarrow$ smECE $\downarrow$
0.1 2.03 1.97 0.26 1.98
0.3–0.8 2.30 2.30 0.25 2.20
0.9 1.43 1.51 0.25 1.44
1.0 2.36 2.34 0.25 2.24
Takeaways.
@@
(2) With CalAttn, performance depends more on $\lambda$ but remains stable near the default. The best calibration in this sweep occurs at $\lambda=0.9$; however, $\lambda=0.1$ remains within $0.6$ percentage points in ECE, making it a reasonable default that avoids dataset-specific tuning.
Appendix J Additional Ablations: ViT-based SATS
J.1 Experimental Protocol
Backbones and datasets.
We use the same ViT backbones and training recipes as in the main paper and report results on CIFAR-10 and CIFAR-100. Unless otherwise stated, all models are trained for 350 epochs with identical data augmentation, optimisers, and learning-rate schedules.
Additional baselines.
Our focus is a ranking-preserving, lightweight scalar calibrator. More expressive post-hoc alternatives (e.g., classwise or vector scaling) can correct class-conditional biases at the expense of additional degrees of freedom and potential decision changes. We include an extended-head comparison (e.g., Dirichlet-style parameterization) to illustrate this capacity–complexity trade-off and leave broader calibrator families and ensembles as future work.
Calibration variants.
We compare the following methods: (i) ViT (baseline) trained with cross-entropy; (ii) ViT+CalAttn (our representation-conditioned temperature head with Brier penalty); (iii) ViT+CalAttn+TS (post-hoc global temperature scaling on CalAttn); (iv) ViT+CalAttn+SATS, where SATS is applied post-hoc on top of CalAttn. SATS follows Joy et al. [2023], using a lightweight MLP that predicts a per-sample temperature from logits (and optionally the CLS embedding).
Temperature scaling.
Global TS selects $T\in\{0.1,0.2,\dots,10.0\}$ on a held-out $5\%$ validation split to minimise ECE, following Mukhoti et al. [2020a] for comparability with focal-family baselines.
SATS training.
SATS is trained post-hoc on the same validation split for 10 epochs with early stopping on validation ECE. Unless stated otherwise, the input consists only of pre-softmax logits, matching Joy et al. [2023].
Metrics.
@@
J.2 Results
Across both datasets, CalAttn consistently reduces ECE and smECE relative to the ViT baseline without affecting Top-1 accuracy. Adding a global TS step on top of CalAttn yields the strongest improvements in all calibration metrics. SATS-only improves calibration over the baseline, but underperforms CalAttn+TS, indicating that learning an representation-conditioned temperature during training provides a stronger inductive bias than purely post-hoc fitting. This suggests that calibrating representations during training induces a smoother uncertainty geometry than fitting a temperature regressor post-hoc, which may overfit the validation split.
@@
CIFAR-100 CIFAR-10
Method Top-1 NLL Brier ECE smECE Top-1 NLL Brier ECE smECE
Baseline $66.01\!\pm\!0.22$ $2.608\!\pm\!0.012$ $0.791\!\pm\!0.004$ $8.54\!\pm\!0.26$ $8.63\!\pm\!0.24$ $77.20\!\pm\!0.31$ $0.850\!\pm\!0.020$ $0.224\!\pm\!0.004$ $3.35\!\pm\!0.20$ $3.52\!\pm\!0.22$
CalAttn $66.20\!\pm\!0.35$ $2.514\!\pm\!0.014$ $0.779\!\pm\!0.004$ $6.39\!\pm\!0.21$ $6.48\!\pm\!0.23$ $77.50\!\pm\!0.20$ $0.842\!\pm\!0.018$ $0.210\!\pm\!0.003$ $10\!\pm\!0.15 2.2$$25\!\pm\!0.16$
CalAttn+TS $67.20\!\pm\!0.26$ $\mathbf{2.489\!\pm\!0.011}$ $\mathbf{0.774\!\pm\!0.003}$ $\mathbf{1.42\!\pm\!0.12}$ $\mathbf{1.10\!\pm\!0.14}$ $78.10\!\pm\!0.25$ $\mathbf{0.792\!\pm\!0.017}$ $\mathbf{0.202\!\pm\!0.003}$ $\mathbf{0.92\!\pm\!0.08}$ $\mathbf{1.01\!\pm\!0.09}$
CalAttn+SATS $66.80\!\pm\!0.28$ $2.519\!\pm\!0.015$ $0.780\!\pm\!0.004$ $2.70\!\pm\!0.15$ $3.44\!\pm\!0.20$ $77.70\!\pm\!0.34$ $0.812\!\pm\!0.019$ $0.206\!\pm\!0.003$ $1.12\!\pm\!0.10$ $1.20\!\pm\!0.11$
Table 17: $\downarrow$ AdaECE (%) before and after applying global temperature scaling. We select $T^{\star}$ by minimizing validation ECE with 15 equal-width bins; parentheses report $T^{\star}$. AdaECE is computed with 15 adaptive (equal-count) bins.
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + $\lambda$ BS ($\lambda=0.1$)
Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T)
@@
Table 18: $\downarrow$ Classwise-ECE (%) before and after applying global temperature scaling (15 bins).
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + $\lambda$ BS ($\lambda=0.1$)
Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T)